# MEDIC: Zero-shot Music Editing with Disentangled Inversion Control

### v3

@@ 3.2 Motivation
Figure 1 and Preliminaries in Appendix A reveal that while techniques like DDIM inversion offer an editable base, they fall short of precision. This potentially compromising the essential content preservation. The implementation of Classifier-free Guidance (CFG) further amplifies the accumulated errors.
@@
In the landscape of prompt-based editing (Dong et al., 2023; Kim et al., 2022; Feng et al., 2023), the ability to grasp the subtleties of linguistics and enable more granular cross-modal interactions stands as a formidable challenge.  Hertz et al. (2022) acknowledges that in image editing, the fusion between text and visual modalities happens within the parameterized noise prediction network $\epsilon_{\theta}$. This leads to the development of various attention control techniques that guide the target denoiser network $\hat{\epsilon}_{\theta}$ in the image domain to better align with target prompts. Yet, analogous control mechanisms for non-rigid music editing are noticeably limited.
@@
Taking these insights forward, we introduce Disentangled Inversion Control (DIC), a novel approach to achieve both rigid and non-rigid music editing. DIC strategically disentangles the diffusion process as triple branches, allowing each branch to optimize its functionality. At the same time, the strategy leverages harmonized attention control to facilitate targeted editing, thus aligning with the dual objectives of preserving the original audio essence and ensuring edit relevance. We will first introduce Harmonized Attention Control in Section 3.3 and discuss Disentangled Inversion in Section 3.4.
3.3 Harmonized Attention Control Framework
The denoising architecture denoted as $\epsilon_{\theta}$, is structured as a sequence of fundamental blocks, each comprising a residual block (He et al., 2015) followed by self-attention and cross-attention modules (Vaswani et al., 2023; Dosovitskiy et al., 2020; Liu et al., 2023b). At the denoising step $t$, the output of the $(l-1)$-th block is passed through self-attention and then aligned with textual cues from prompt $P$ within the cross-attention layer. The attention mechanism is formalized as:
$\text{Attention}(Q,K,V)=MV=\text{Softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right)V$ (1)
@@ 3.3 Harmonized Attention Control Framework
where $Q$ denotes the query features derived from the audio features, $K$ and $V$ represent the key and value features, and $M$ is the attention map. We explore varying semantic transformations of audio content through harmonized attention control strategies — cross-attention control for rigid changes and mutual self-attention control for non-rigid adjustments. Finally, we introduce an intermediate branch to host the desired harmonic and melodic information in the target music. The framework of harmonized attention control is depicted in Figure 2.
@@
Figure 2: The framework of Harmonized Attention Control (HAC). HAC unifies cross-attention control and mutual self-attention control with an additional branch named Harmonic Branch to host the desired composition and structural information in the target music.
3.3.1 Cross-Attention Control
@@
Cross-attention Control (CAC) aims to inject the attention maps that are obtained from the generation with the original prompt ${\mathcal{P}}$, into a second generation with the target prompt ${\mathcal{P}}^{*}$. We follow the practice of Prompt-to-Prompt (Hertz et al., 2022) and define CAC as Global Attention Refinement and Local Attention Blend.
@@
At a given time step $t$, the attention map $M_{t}$ for both the origin and target branches is computed, averaging over all layers with respect to the noised latent $\mathbf{z}_{t}$. We employ an alignment function $A$ that maps each token index from the target prompt $\mathcal{P}^{*}$ to its equivalent in $\mathcal{P}$ or to None for non-aligning tokens. The refinement action is thus:
$\text{Refine}(M^{src}_{t},M^{tgt}_{t},t)=\begin{cases}(M^{tgt}_{t})_{i,j}&\text{if }A(j)=\text{None},\\ (M_{t})_{i,j}&\text{otherwise}.\end{cases}$ (2)
@@
Beyond global attention, we incorporate a blending mechanism as suggested by  Hertz et al. (2022) and  Mokady et al. (2023). This method selectively integrates and maintains certain semantics by using target blend words $w^{tgt}$ for semantic additions and source blend words $w^{src}$ for semantic preservation. At each denoising step $t$, the mechanism operates on the target latent ${\bm{z}}^{tgt}_{t}$ as follows:
$m_{\text{tgt}}$ $=\text{Threshold}\left[M^{tgt}_{t},w_{\text{tgt}},k_{\text{tgt}}\right],$ (3)
$m_{\text{src}}$ $=\text{Threshold}\left[M^{src}_{t},w_{\text{src}},k_{\text{src}}\right],$ (4)
${\bm{z}}^{tgt}_{t}$ $=(1-m^{tgt}+m^{src})\odot{\bm{z}}^{src}_{t}+(m^{tgt}-m^{src})\odot{\bm{z}}^{tgt}_{t}$ (5)
where $m_{\text{tgt}}$ and $m_{\text{src}}$ are binary masks and threshold function is as delineated below:
$\text{Threshold}(M,k)=\begin{cases}1&\text{if }M_{i,j}\geq k,\\ 0&\text{if }M_{i,j}<k.\end{cases}$ (6)
@@
${\bm{z}}^{tgt}_{t}=\text{LocalEdit}({\bm{z}}^{src}_{t},{\bm{z}}^{tgt}_{t},M^{src}_{t},M^{tgt}_{t},w_{\text{src}},w_{\text{tgt}})$ (7)
@@
Applying cross-attention control only at early stages ensures creative flexibility while maintaining structure. Acting on insights from  Hertz et al. (2022), we limit cross-attention to the initial phases up to a cutoff point $\tau_{c}$. This moderation allows us to capture the nuances and intended changes in musical compositions effectively. The approach is defined as:
$\text{CrossEdit}(M^{src},M^{tgt},t)=\begin{cases}\text{Refine}(M^{src}_{t},M^{tgt}_{t})&\text{if}\ t\geq\tau_{c},\\ M^{tgt}_{t}&\text{if}\ t<\tau_{c}.\end{cases}$ (8)
3.3.2 Mutual Self-Attention Control
We diverge from the conventional use of cross-attention mechanisms and instead draw inspiration from the MasaCtrl (Cao et al., 2023) technique to refine music structure through self-attention queries. These queries adeptly navigate through non-rigid musical transformations, aligning with the designated musical theme or instrument (target prompt). The process begins by sketching the foundational musical theme using the target’s self-attention components—$Q^{tgt}$, $K^{tgt}$, and $V^{tgt}$. This is followed by enriching this theme with elements resembling the thematic content from the source ($K^{src}$, $V^{src}$), steered by $Q^{tgt}$. However, applying this attentive modulation uniformly over all processing layers and through every denoising step might result in a composition excessively mirroring the source. Consequently, echoing the ethos of MasaCtrl, our proposed solution selectively employs mutual self-attention in the decoder portion of our music editing U-Net, initiated after a set number of denoising iterations.
@@ 3.3.2 Mutual Self-Attention Control
The application of mutual self-attention is meticulously planned, beginning at a specific denoising step $S$ and extending beyond a designated layer $L$. The strength and influence of this control mechanism are designed as follows:
$\text{SelfEdit}({Q^{src},K^{src},V^{src}},{Q^{tgt},K^{tgt},V^{tgt}},t)=\begin{cases}{Q^{src},K^{src},V^{src}}&\text{if }t\geq S\text{ and }l\geq L,\\ {Q^{tgt},K^{src},V^{src}}&\text{otherwise}\\ \end{cases}$ (9)
In this framework, $S$ signifies the denoising step from which the mutual self-attention control commences, serving as a temporal threshold. Similarly, $L$ distinguishes the layer index below which this nuanced control strategy becomes operational, tailoring the musical output towards the intended artistic direction.
3.3.3 Harmonic Branch Integretion
The naive combination of cross-attention control and mutual self-attention control sequentially would lead to sub-optimal results in the original dual-branch setup, especially failing the global attention refinement. To address this issue, we introduce an additional latent harmonic branch, which serves as an intermediate to host the desired composition and structural information in the target music.
Our unified framework is detailed in Algorithm 2. During each forward step of the diffusion process, we start with mutual self-attention control on $z^{src}$ and $z^{tgt}$ and assign the output to the harmonic branch latent $z^{har}$. This lays the formal structure of the target music. Following this, cross-attention control is applied on $M^{har}$ and $M^{tgt}$ to refine the semantic information for $M^{tgt}$. As illustrated in Figure 2, the harmonic branch output $z^{har}_{0}$ reflects the requested non-rigid changes (e.g., “violin"), while preserving the rigid content semantics (e.g., “with noise"). The target branch output $z^{tgt}_{0}$ builds upon the structural layout of the $z^{har}$ while reflecting the requested rigid changes (e.g., “with noise").
@@ 3.3.3 Harmonic Branch Integretion
2:  $\epsilon_{\text{src}},\{Q^{\text{src}},K^{\text{src}},V^{\text{src}}\},M_{\text{src}}=\epsilon_{\theta}(z^{\text{src}}_{\tau},\tau,c_{\text{src}})$
3:  $\epsilon_{\text{tgt}},\{Q^{\text{tgt}},K^{\text{tgt}},V^{\text{tgt}}\},M_{\text{tgt}}=\epsilon_{\theta}(z^{\text{tgt}}_{\tau},\tau,c_{\text{tgt}})$
4:  $\{Q^{\text{har}},K^{\text{har}},V^{\text{har}}\}=\text{SelfEdit}(\{Q^{\text{src}},K^{\text{src}},V^{\text{src}}\},\{Q^{\text{tgt}},K^{\text{tgt}},V^{\text{tgt}}\},\tau)$
5:  $\epsilon_{\text{har}},M_{\text{har}}=\epsilon_{\theta}(z^{\text{har}}_{\tau},\tau,c_{\text{src}};\{Q^{\text{har}},K^{\text{har}},V^{\text{har}}\})$
6:  $\hat{M^{\text{tgt}}}=\text{CrossEdit}(M_{\text{har}},M_{\text{tgt}},\tau)$
7:  $\hat{\epsilon_{\text{tgt}}}=\epsilon_{\theta}(z^{\text{tgt}}_{\tau},\tau,c_{\text{tgt}};\hat{M^{\text{tgt}}})$
8:  $z^{\text{src}}_{\tau-1},z^{\text{tgt}}_{\tau-1},z^{\text{har}}_{\tau-1}=\text{Sample}([z^{\text{src}}_{\tau},z^{\text{tgt}}_{\tau},z^{\text{har}}_{\tau}],[\epsilon_{\text{src}},\epsilon_{c_{\text{tgt}}},\epsilon_{\text{har}}],\tau)$
9:  ${\bm{z}}^{tgt}_{\tau-1}$ = $\text{LocalEdit}({\bm{z}}^{src}_{\tau-1},{\bm{z}}^{tgt}_{\tau-1},M^{src}_{\tau-1},M^{tgt}_{t},w_{src},w_{tgt})$
@@
3.4 Disentangled Inversion Technique
Recognizing the limitations of using DDIM inversion without classifier-free guidance, we observe that it yields an easily modifiable but imprecise approximation of the original audio signal. Increasing the guidance scale enhances editability, but sacrifices reconstruction accuracy due to latent code deviation during editing.
In order to address this issue, our methodology, which we have termed the Disentangled Inversion Technique disentangles into three branches: the source, the harmonic, and the target branch. This decoupling is designed to unleash the capabilities of each branch separately. For the source branch, we implement a targeted correction mechanism. By reintegrating the distance $z^{*}_{t}-z^{src}_{t}$ into $z^{src}_{t}$, we directly mitigate the deviation of the pathway. This straightforward adjustment effectively rectifies the path and minimizes the accumulated errors introduced by both DDIM inversion and classifier-free guidance, thereby enhancing consistency in the reconstructed audio. On the other hand, the target branch and harmonic branch are left unmodified to fully leverage the innate capabilities of diffusion models in generating the desired target audio. The branches’ untouched state ensures that the model’s potential is utilized to its fullest extent, thereby ensuring the fidelity and integrity of the generated audio. We will further discuss this in the Section 4.2.2. The algorithm of Disentangled Inversion Technique has been outlined in Algorithm 2.
Typical diffusion-based editing (Han et al., 2023b; Miyake et al., 2023) involves two parts: an inversion process to get the diffusion space of the audio, and a forward process to perform editing on the diffusion space. Disentangled Inversion can be plug-and-played into the forward process and rectifies the deviation path step by step. Specifically, Disentangled Inversion first computes the difference between ${\bm{z}}^{*}_{t-1}$ and ${\bm{z}}^{src}_{t-1}$, then adds the difference back to ${\bm{z}}^{src}_{t-1}$ in DDIM forward. We only add the difference of the source prompt in latent space and update ${\bm{z}}^{src}_{t-1}$, which is the key to retaining the editability of the target prompt’s latent space.
@@ 3.4 Disentangled Inversion Technique
5:  for $t=T$ to $1$ do
6:     $[{\bm{d}}^{src}_{t-1},{\bm{d}}^{tgt}_{t-1},{\bm{d}}^{har}_{t-1}]\leftarrow{\bm{z}}^{*}_{t-1}$ DDIM_Forward$({\bm{z}}^{src}_{t},t,[{\mathcal{P}},{\mathcal{P}}^{*},{\mathcal{P}}],\omega)$
7:     ${\bm{z}}^{src}_{t-1}\leftarrow$ DDIM_Forward(${\bm{z}}^{src}_{t},t,[{\mathcal{P}},{\mathcal{P}}^{*},{\mathcal{P}}],\omega$) + $[{\bm{d}}^{src}_{t-1},{\bm{0}},{\bm{0}}]$
8:     ${\bm{z}}^{har}_{t-1}\leftarrow$ DDIM_Forward(${\bm{z}}^{har}_{t},t,[{\mathcal{P}},{\mathcal{P}}^{*},{\mathcal{P}}],\omega$) + $[{\bm{d}}^{src}_{t-1},{\bm{0}},{\bm{0}}]$
9:     ${\bm{z}}^{tgt}_{t-1}\leftarrow$ DDIM_Forward(${\bm{z}}^{tgt}_{t},t,[{\mathcal{P}},{\mathcal{P}}^{*},{\mathcal{P}}],\omega$) + $[{\bm{d}}^{src}_{t-1},{\bm{0}},{\bm{0}}]$
@@
11:  return  ${\bm{z}}^{tgt}_{0}$
4 Experiments
Implementation Details. We infer different editing methods using the pre-trained AudioLDM 2 (Liu et al., 2023c) models with 200 inference steps. The setting of baselines is followed by Manor & Michaeli (2024) and is all evaluated in NVIDIA A800 GPU for a fair comparison. We evaluate all methods in ZoMo-Bench across all editing types with fixed length. And we use the MusicDelta subset of the MelodyDB (Bittner et al., 2014) dataset for variable length comparisons, comprised of 34 musical excerpts in varying styles and in lengths ranging from 20 seconds to 5 minutes. Details can be found in Appendix C.
Evaluation Metrics. Our models employ a comprehensive evaluation using both objective and subjective metrics to assess essential content preservation, text-audio alignment fidelity, and audio quality. Objective metrics include Structure Distance (SD) (Ju et al., 2023), CLAP (Contrastive Language-Audio Pretraining) Score (Elizalde et al., 2023), LPAPS (Learned Perceptual Audio Patch Similarity) (Iashin & Rahtu, 2021; Paissan et al., 2023), and FAD (Fréchet Audio Distance) (Kilgour et al., 2018). The CLAP Score evaluates how well the edited audio aligns with the target prompt, while Structure Distance, LPAPS, and FAD, which have been adapted from image domain metrics, measure the similarity between the edited and source audio. Specifically, LPAPS quantifies the consistency of the edited audio in relation to the source audio, and the FAD metric assesses the distance between two distributions of audio signals. The Structure Distance measures the structural similarity between the edited and source audio. For the subjective evaluation, we conduct crowd-sourced human assessments using the Mean Opinion Score (MOS) to evaluate both editing fidelity (MOS-Q) and content preservation (MOS-P). We attach the details of all metrics in Appendix C.1.
4.1 Zero-shot Music Editing Results

### v4

@@ 4.2. Disentangled Inversion Control
Preliminaries and Figure 4 reveal that while techniques like DDIM inversion offer an editable base, they fall short of precision, potentially compromising essential content preservation. The implementation of Classifier-free Guidance (CFG) further amplifies the accumulated errors.
@@
In the landscape of prompt-based editing (dong2023prompt; kim2022diffusionclip; feng2023trainingfree), grasping linguistic subtleties and enabling more granular cross-modal interactions remains a formidable challenge.  hertz2022prompt acknowledges that in image editing, the fusion between text and visual modalities occurs within the parameterized noise prediction network $\epsilon_{\theta}$, which leads to the development of various attention control techniques that guide the target denoiser network $\hat{\epsilon}_{\theta}$ in the image domain to better align with target prompts. However, similar control mechanisms for non-rigid music editing are noticeably limited.
@@
Taking these insights forward, we introduce Disentangled Inversion Control (DIC), a novel approach to achieve both rigid and non-rigid music editing. DIC strategically disentangles the diffusion process as triple branches (i.e. source branch, harmonic branch, and target branch), and allows each branch to optimize its functionality. At the same time, the DIC strategy leverages our proposed harmonized attention control to facilitate targeted editing, thus aligning with the dual objectives of preserving the original audio essence and ensuring edit fidelity. Next, we first introduce Harmonized Attention Control in Section 4.3 and discuss Disentangled Inversion in Section 4.4.
4.3. Harmonized Attention Control Framework
@@
The denoising architecture $\epsilon_{\theta}$ is structured as a sequence of fundamental blocks, each comprising a residual block (he2015deep) followed by self-attention and cross-attention modules (vaswani2023attention; dosovitskiy2020image; liu2023vit). The proposed harmonized attention control (HAC) framework is depicted in Figure 5. We explore varying semantic transformations of audio content through harmonized attention control strategies - cross-attention control for rigid editings and mutual self-attention control for non-rigid editings. We introduce an intermediate Harmonic Branch to host the desired harmonic and melodic information in the target music. Below we elaborate on Cross-Attention Control, Mutual Self-Attention Control, and Harmonic Branch Integration within HAC.
@@
Figure 5. The Harmonized Attention Control (HAC) framework. HAC unifies cross-attention control and mutual self-attention control with an additional branch named Harmonic Branch to host the desired composition and structural information in the target music.
4.3.1. Cross-Attention Control
@@
Cross-attention Control (CAC) aims to inject the attention maps that are obtained from the generation with the original prompt $\mathcal{P}$}, into a second generation with the target prompt $\mathcal{P}$$^{*}$. Motivated by Prompt-to-Prompt (hertz2022prompt), We define CAC as Global Attention Refinement and Local Attention Blend.
@@
At a given time step $t$, the attention map $M_{t}$ for both source and target branches is computed, averaging over all layers with respect to the noised latent $\mathbf{z}_{t}$. We employ an alignment function $A$ that maps each token index from the target prompt $\mathcal{P}^{*}$ to its equivalent in $\mathcal{P}$, or to None for unaligned tokens. The refinement action is:
(6) $\text{Refine}(M^{src}_{t},M^{tgt}_{t},t)=\begin{cases}(M^{tgt}_{t})_{i,j}&\text{if }A(j)=\text{None},\\ (M_{t})_{i,A(j)}&\text{otherwise}.\end{cases}$
@@
Beyond global attention, we incorporate a blending mechanism suggested by hertz2022prompt and  mokady2023null. This method selectively integrates and maintains certain semantics by using target blend words $w^{tgt}$, which are words in the target prompt whose semantics need to be added; and source blend words $w^{src}$, which are words in the source prompt whose semantics need to be preserved. At each denoising step $t$, the mechanism operates on the target latent ${\bm{z}}^{tgt}_{t}$ as follows:
(7) $m_{\text{tgt}}$ $=\text{Threshold}\left[M^{tgt}_{t}(w_{\text{tgt}}),k_{\text{tgt}}\right],$
(8) $m_{\text{src}}$ $=\text{Threshold}\left[M^{src}_{t}(w_{\text{src}}),k_{\text{src}}\right],$
(9) ${\bm{z}}^{tgt}_{t}$ $=(1-m^{tgt}+m^{src})\odot{\bm{z}}^{src}_{t}+(m^{tgt}-m^{src})\odot{\bm{z}}^{tgt}_{t}$
where $m_{\text{src}}$ and $m_{\text{tgt}}$ are binary masks obtained by calibrating the aggregated attention maps $M^{src}_{t}(w^{src})$ and $M^{tgt}_{t}(w^{tgt})$ with threshold parameters $k_{src}$ and $k_{tgt}$, using the following threshold function:
(10) $\text{Threshold}(M,k)=\begin{cases}1&\text{if }M_{i,j}\geq k,\\ 0&\text{if }M_{i,j}<k.\end{cases}$
@@
(11) ${\bm{z}}^{tgt}_{t}=\text{LocalEdit}({\bm{z}}^{src}_{t},{\bm{z}}^{tgt}_{t},M^{src}_{t},M^{tgt}_{t},w_{\text{src}},w_{\text{tgt}})$
@@
Implementing cross-attention control throughout the entire sampling schedule can cause excessive focus on music structures, hindering the ability to incorporate intended changes. In contrast, applying it only during the early stages allows for creative flexibility while still preserving structural integrity. Therefore, we limit cross-attention to the initial phases up to a cutoff point $\tau_{c}$. This moderation allows us to effectively capture the nuances and intended changes in musical compositions. The scheduling approach is defined as follows:
(12) $\text{CrossEdit}(M^{src},M^{tgt},t)=$
[…]
4.3.2. Mutual Self-Attention Control
[…]
@@ 4.3.2. Mutual Self-Attention Control
The application of the nuanced mutual self-attention control is meticulously planned, beginning at a specific denoising step $S$ and extending beyond a designated layer index $L$. The strength and influence of this control mechanism are designed as follows.
(13) $\text{SelfEdit}({Q^{src},K^{src},V^{src}},{Q^{tgt},K^{tgt},V^{tgt}},t)=$
[…]
In this framework, $S$ serves as a temporal threshold while $L$ tailors the musical output towards the intended artistic direction.
4.3.3. Harmonic Branch Integretion
We hypothesize that a simple sequential combination of cross-attention control and mutual self-attention control may yield sub-optimal results within the original dual-branch framework. This approach is particularly ineffective in refining global attention, as depicted in Figure 4(d). Our empirical validation, presented in Table 2 (HAC vs w/o HB), supports this hypothesis. To address this issue, we introduce an additional latent harmonic branch, which serves as an intermediate branch to host the desired composition and structural information in the target music. The unified framework of Harmonized Attention Control is detailed in Algorithm 1. During each forward step of the diffusion process, we start with mutual self-attention control on $z^{src}$ and $z^{tgt}$ and assign the output to the harmonic branch latent $z^{har}$. This latent lays the formal structure of the target music. Next, cross-attention control is applied on $M^{har}$ and $M^{tgt}$ to refine the semantic information for $M^{tgt}$. As illustrated in Figure 5, the harmonic branch output $z^{har}_{0}$ reflects the requested non-rigid changes (e.g., “rapid”), while preserving the rigid content semantics (e.g., “guitar”). The target branch output $z^{tgt}_{0}$ builds upon the structural layout of the $z^{har}$ while reflecting the requested rigid changes (e.g., “violin”).
@@ 4.3.3. Harmonic Branch Integretion
[…]
@@
4.4. Disentangled Inversion Technique
Using DDIM inversion without classifier-free guidance yields an easily modifiable but imprecise approximation of the original audio signal. Increasing the classifier-free guidance scale enhances editability, but sacrifices reconstruction accuracy due to latent code deviation during editing.
In order to address these limitations, we propose Disentangled Inversion Technique to disentangle the diffusion process into three branches: the source branch, the harmonic branch, and the target branch, with the detailed algorithm outlined in Algorithm 2. This decoupling is designed to unleash the capabilities of each branch separately. For the source branch, we implement a targeted correction mechanism. By reintegrating the distance $z^{*}_{t}-z^{src}_{t}$ into $z^{src}_{t}$, we directly mitigate the deviation of the pathway. This straightforward adjustment effectively rectifies the path and minimizes the accumulated errors introduced by both DDIM inversion and classifier-free guidance, thereby enhancing consistency in the reconstructed audio. On the other hand, the target branch and harmonic branch are left unmodified to fully leverage the innate capabilities of diffusion models in generating the desired target audio, thereby ensuring the fidelity and integrity of the generated audio. Effectiveness of Disentangled Inversion Technique are verified and discussed in Section 2.
Typical diffusion-based editing (han2023improving; miyake2023negativeprompt) involves two parts: an inversion process to obtain the diffusion space of the audio, and a forward process to perform editing on the diffusion space. In contrast, our Disentangled Inversion Technique can be plug-and-played into the forward process and rectifies the deviation path step by step. Specifically, Disentangled Inversion first computes the difference between ${\bm{z}}^{*}_{t-1}$ and ${\bm{z}}^{src}_{t-1}$, then adds the difference back to ${\bm{z}}^{src}_{t-1}$ in DDIM forward. We only add the difference of the source prompt in latent space and update ${\bm{z}}^{src}_{t-1}$, which is the key to retaining the editability of the latent space of the target prompt.
@@ 4.4. Disentangled Inversion Technique
5:for $t=T$ to 11 do
6:  ${\bm{d}}^{src}_{t-1},{\bm{d}}^{tgt}_{t-1},{\bm{d}}^{har}_{t-1}]\leftarrow{\bm{z}}^{*}_{t-1}$
[…]
7:  ${\bm{z}}^{src}_{t-1}\leftarrow$ DDIM_Forward(${\bm{z}}^{src}_{t},t,[{\mathcal{P}},{\mathcal{P}}^{*},{\mathcal{P}}],\omega$) + ${\bm{d}}^{src}_{t-1},{\bm{0}},{\bm{0}}]$
8:  ${\bm{z}}^{har}_{t-1}\leftarrow$ DDIM_Forward(${\bm{z}}^{har}_{t},t,[{\mathcal{P}},{\mathcal{P}}^{*},{\mathcal{P}}],\omega$) + ${\bm{d}}^{src}_{t-1},{\bm{0}},{\bm{0}}]$
9:  ${\bm{z}}^{tgt}_{t-1}\leftarrow$ DDIM_Forward(${\bm{z}}^{tgt}_{t},t,[{\mathcal{P}},{\mathcal{P}}^{*},{\mathcal{P}}],\omega$) + ${\bm{d}}^{src}_{t-1},{\bm{0}},{\bm{0}}]$
@@
11:Return: ${\bm{z}}^{tgt}_{0}$
5. Experiments
Table 1. Performance comparison of objective and subjective zero-shot music editing on ZoME-Bench (fixed length) and MusicDelta (variable length). The evaluation focuses on content preservation (SD, LPAPS, FAD, Chroma, MOS-P) and edit fidelity (CLAP Score, Accuracy, MOS-Q). Results include mean and standard deviation for both edit fidelity (MOS-Q) and content preservation (MOS-P). Metrics achieving the best scores are highlighted in bold.